OpenStudy (anonymous):

Calculus question: Find the indefinite integral using substitution and change of variables.

OpenStudy (anonymous):

\[\int\limits_{}^{}x \]\[\sqrt{x+6}dx\]

OpenStudy (anonymous):

^ The above problem should be together. Could someone walk me through this problem?

OpenStudy (zarkon):

this \[\int x\sqrt{x+6}\,\,dx\] ?

OpenStudy (anonymous):

yes, when i tried to put the square root sign in, it went to the front of the integral sign. Thanks.

OpenStudy (zarkon):

let \[u=x+6\] \[du=dx\] and \[x=u-6\] \[\int x\sqrt{x+6}\,\,dx=\int (u-6)\sqrt{u}\,\,du=\cdots\]

OpenStudy (zarkon):

you need more help than that?

OpenStudy (anonymous):

I'll let you know, I'm gonna see if I can finish it from there

OpenStudy (anonymous):

In my book before it simplifies the integral to get the final answer it gets this: \[\int\limits_{}2/5 (x+6)^{5/2}-4(x+6)^{3/2}+C\] I don't get how they got -4(x+6)^3/2

OpenStudy (zarkon):

\[\int (u-6)\sqrt{u}\,\,du=\int(u^{3/2}-6u^{1/2})du\] \[=\frac{2}{5}u^{5/2}-6\frac{2}{3}u^{3/2}+c=\frac{2}{5}u^{5/2}-4u^{3/2}+c\] \[=\frac{2}{5}(x+6)^{5/2}-4(x+6)^{3/2}+c\]

OpenStudy (anonymous):

oh okay, I see now. Thanks .